The following reaction

2H2S(g)⇌2H2(g)+S2(g),Kc=1.654×10−7 at 800∘C

is carried out at the same temperature with the following initial concentrations: [H2S]=0.132M, [H2]=0.165 M, and [S2]=0.00 M. Find the equilibrium concentration of S2.

Express the molarity to three significant figures. Answer in units of nM.
Answer #1

Answer:

Given reaction is

2H2S(g)⇌2H2(g)+S2(g), Kc=1.654×10−7

Initial concentrations: [H2S]=0.132M, [H2]=0.165 M, and [S2]=0.00 M

2H2S(g) ⇌ 2H2(g)+S2(g)

Initial 0.132 0.165 0

Change -2x +2x +x

Equilibrium 0.132-2x 0.165+2x x

Kc=[H2]2[S2]/H2S]2

(1.654×10−7) = (0.165+2x)2(x)/(0.132-2x)2

Since Kc is very small then 0.165+2x ~ 0.165 and 0.132-2x ~ 0.132.

Therefore (1.654×10−7) (0.132)2/(0.165)2=x

x=1.058×10−7

Therefore the equilibrium concentration of S2, [S2]=x=1.058×10−7 M

Since 1nM= 1 x 10-9 M

The equilibrium concentration of S2, [S2]=105.8 nM ~ 106 nM (3 significant figures)

Please let me know if you have any doubt. Thanks.